Solve the following equation by using the quadratic formula. (Enter solutions from smallest to largest. If there is one solution with a multiplicity of two, enter it in both answer boxes.)
(2x-1)(x+2)=9
x = ?
x = ?
Please show me the steps to get the answer. Thanks

Answer by rfer(16322)

(2x-1)(x+2)=
2x^2+4x-x-2=
2x^2+3x-2=9
2x^2+3x-11=0
a=2, b=3, c=-11
(-3+-\/9-4*2*-11)/4
(-3+-\/97)/4
(-3-9.85)/4=-3.2125 (-3.2125,0)
(-3+9.85)/4=+1.7125 (+1.7125,0)